OpenStudy (anonymous):

if x^2+bx+c= (x+e) (x+f) and c is postive and b is negative, what can you say about the sighs of e and f?

OpenStudy (anonymous):

e and f are negative

OpenStudy (anonymous):

negative

OpenStudy (anonymous):

i thought it was postive

OpenStudy (anonymous):

just try it with some examples

OpenStudy (anonymous):

see the sum of roots of a quadratic is -b/a a=1 here if b is -ve then the sum of roots is positive the product is c/a, and c is +ve, so product is +ve, which means their sign is same so both roots are positive now the roots of (x+e)(x+f) are -e and -f, and if these are positive then e and f need to be negative, as - x - = +

OpenStudy (amistre64):

x^2 -(e+f)x + (ef) ef = + number (+e)(+f) OR (-e)(-f) e+f = - number (+e)+(+f) = + number; +1+2 = +3 (-e)+(-f) = - number; (-1)+(-2) = -3

OpenStudy (amistre64):

ax^2 (+-) bx (+-) c ^ ^ ^ tells us to add or subtract ^ tells us the bigger factors sign

OpenStudy (amistre64):

x^2 -4x -12 ^ ^ subtract factors to get 7 ^ the bigger factor is (-) 6(2) = 12 -6+2 = -4 (x-6) (x+2)

OpenStudy (amistre64):

umm .... subtract to get 4 .....

OpenStudy (anonymous):

so i get a postive?